OpenStudy (brooke..help00):

Part A: Solve −np − 90 < 30 for n. Show your work. Part B: Solve 5m − 3h = 45 for h. Show your work.

RhondaSommer (rhondasommer):

Part A:\[-np<120\\n>\frac{120}{-p}\] Part B:\[-3h=-5m+45\\h=\frac{5}3m+45\]